OpenStudy (anonymous):

Factor and find roots. 6x^2+x-1

OpenStudy (anonymous):

Find two factors of 1 -6 that add up to 1.

OpenStudy (anonymous):

Which would be..?

OpenStudy (anonymous):

If you can do that, I can help you to your answer...

OpenStudy (anonymous):

Finding those would be factoring it and therefore my answer. My problem is that I'm unable to find them.

OpenStudy (anonymous):

I mean factors of -6 that add up to 1.

OpenStudy (anonymous):

You wouldn't quite have your answer after that.

OpenStudy (anonymous):

Oh, 3 and -2?

OpenStudy (anonymous):

Yes. Then split up the middle term using that. \[6x^2 + 3x - 2x - 1 \implies (6x^2 + 3x) + (-2x - 2)\]Can you factor out the GCF in each of those groupings?

OpenStudy (anonymous):

\[(2x^2+x) and (-x-1)\]

OpenStudy (anonymous):

Not quite. \[3x(2x + 1) - (2x + 1)\]Now you can use the distributive property and combine them :) \[(3x - 1)(2x + 1)\]If it doesn't match up, it's because I had a typo above... There is you're final answer though :)

OpenStudy (anonymous):

So how does the 3x come into play when you use the distributive?

OpenStudy (anonymous):

That's what you factor out. \[(6x^2 + 3x) + (-2x - 1)\]See how you can factor out a 3x and factor out a -1?

OpenStudy (anonymous):

So what are the 2 zeros?

OpenStudy (anonymous):

The original question asks to factor. Not to find the solutions. If you mean the factors, then it's (3x - 1)(2x + 1)

OpenStudy (anonymous):

I forgot to put that into the original question x~x

OpenStudy (anonymous):

So you're saying you need to find x?

OpenStudy (anonymous):

3x - 1 = 0 3x = 1 x = 1/3 2x + 1 = 0 2x = -1 x = -1/2 x = 1/3, -1/2